Circular No. 8148

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003fe, 2003ff, 2003fg, 2003fh, 2003fi, 2003fj, 2003fk R. Ellis, California Institute of Technology; M. Sullivan, University of

Durham; and P. Nugent, Lawrence Berkeley National Laboratory, report on CCD spectra (range 360–950 nm) of the Legacy Survey supernovae announced on IAUC 8147, obtained during Apr. 2–6 (Keck II + DEIMOS) and on Apr. 7, May 6, June 1, and 2 (Keck I + LRIS). Host-galaxy redshifts were measured using the [O II] 372.7-nm line or the Ca II H and K lines, and the supernova features were consistent with these redshifts. The epochs for the six type-Ia supernovae (all except SN 2003fe) are based on a comparison of the spectroscopic features with the templates described by Nugent et al. (2002, PASP 114, 803). SN 2003fe (Apr. 2) is of type II, with a redshift of 0.42 (at early epoch, with a featureless blue spectrum). Redshifts of the type-Ia supernovae (which were all near peak brightness, with SNe 2003fh and 2003fi observed a week before peak): SN 2003ff, 0.50 (Apr. 3, 6, 7); SN 2003fg, 0.24 (May 6); SN 2003fh, 0.25 (May 6); SN 2003fi, 0.44 (June 1, 2); SN 2003fj, 0.47 (June 2); SN 2003fk, 0.35 (June 1).

SUPERNOVAE 2003fl, 2003fm, 2003fn, 2003fo, 2003fp, 2003fq Further to IAUC 8147, P. Astier et al. report additional apparent su-

pernovae during the CFHT Legacy Survey:

SN  2003 UT  α_{2000}  δ_{2000}  i′  Offset
2003fl Mar. 23.53 14^{h}19^{m}55.^{s}64 +53^{o}00′51′′.9 23.8 0′′.3 W, 0′′.3 N
2003fm Mar. 23.53 14 21 07.10 +52 43 27.5 24.0 1′′.7 W, 0′′.5 N
2003fn Apr. 24.54 14 16 33.44 +52 20 32.1 21.5 0′′.1 W, 1′′.6 N
2003fo Apr. 24.54 14 17 58.43 +53 28 57.5 23.2 0′′.6 E, 0′′.6 S
2003fp Apr. 24.54 14 20 53.61 +52 36 20.7 22.9 0′′.5 E, 0′′.3 N
2003fq May 24.45 14 19 46.84 +52 30 27.7 22.2 0′′.2 W, 0′′.2 N

Additional i′ magnitudes: SN 2003fl, Mar. 3 UT, [24.5; 25.5, 23.6; Apr. 1.5, 22.8. SN 2003fm, Mar. 3, [24.5; 25.5, 23.6; Apr. 1.5; 23.7. SN 2003fn, Mar. 3, [24.5; Apr. 29.5, 21.6. SN 2003fo, Mar. 3, [24.5; Apr. 29.5, 22.4. SN 2003fp, Mar. 3, [24.5; Apr. 29.5, 22.6. SN 2003fq, Mar. 23, [24.5; May 28.4, 22.9; June 2.4, 23.2.

Ellis et al. provide the following supernova types and redshifts from their spectra (obtained as above): SN 2003fl, ?, 0.68 (Apr. 5); SN 2003fm, ?, 0.83 (Apr. 4, 6); SN 2003fn, Ia, 0.29 (May 6; 2 weeks past peak brightness); SN 2003fo, Ia, 0.37 (May 6; 1 week before peak brightness); SN 2003fp, Ia, 0.45 (May 6; near peak brightness).

2003 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green